what is the equation of the line with a slope of -2 and passing through the point (-3,5)
Answer by ikleyn(53937)

.
y - 5 = (-2)*(x-(-3)).


Or, which is the same,


y - 5 = -2*(x+3).

-------------------

If you need more explanations,  or if you want to learn the subject from  "a"  to  "Z"  on your own,  look into the lesson
    - Equation for a straight line having a given slope and passing through a given point
in this site.

My other lessons in this site related to the slope of a straight line are

    - Find the slope of a straight line in a coordinate plane passing through two given points
    - Equation for a straight line having a given slope and passing through a given point
    - Solving problems related to the slope of a straight line
    - Equation for a straight line in a coordinate plane passing through two given points
    - Equation for a straight line parallel to a given line and passing through a given point
    - Equation for a straight line perpendicular to a given line and passing through a given point

    - OVERVIEW of lessons related to the slope of a straight line

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.